# of our esteemed colleague Professor

### v1

@@ 5. Even dimensional case
The following has been obtained in [7, Theorem 1.1] by using cumulative winding number in $\frac{1}{2}\mathbb{Z}$. Here, we give another proof, which seems to be more elementary in nature.
@@
We can construct a generic homotopy between the image simple fold maps $\tilde{w}$ and $\tilde{z}\colon S^{1}\tilde{\times}S^{2n-1}\to\mathcal{M}$ as follows. We first apply the birth move (see [15, Fig. 5]) to $\tilde{w}$ in such a way that we have a wrinkle consisting of two cusps together with fold arcs of indices n1n-1 and $n$ near the center circle of $\mathcal{M}$, i.e., inside the region $R$ bounded by $)\mathcal{M}\setminus\tilde{w}(\Sigma(\tilde{w}$)). Note that if we reverse the normal direction, the indices turn to $n$ and n1n-1. Then, we merge the two cusps using a path connecting them in the source manifold whose image is embedded in $R$. This is possible for $n>1$, since the fibers over the points in $R$ are connected. For $n=1$, the fibers are not connected: however, we can use a similar cusp elimination technique as in [11]. Then, we get an image simple fold map equivalent to $\tilde{z}$. The singular point set of the homotopy thus constructed is the disjoint union of a Möbius band and a cylinder, and it has no triple self-intersections.
@@ Acknowledgment
The authors would like to thank David Gay who kindly explained to the authors how to construct generalized Dehn twists. They are also thankful to Liam Kahmeyer for discussions of the topic. The second author would also like to thank the Department of Mathematics, Kansas State University, for the hospitality during the preparation of the manuscript.

### v2

@@ 6. Even dimensional case
The following has been obtained in [10, Theorem 1.1] by using cumulative winding number in $\frac{1}{2}\mathbb{Z}$. Here, we give another proof, which seems to be more elementary in nature.
@@
Remark 6.3.
In the case of an image simple fold map $f$ of a closed orientable surface $M$ to $\mathbb{R}^{2}$, the components of $\Sigma(f)$ can be signed so that the algebraic number of components of $\Sigma(f)$ is equal to $\chi(M)/2$. Specifically, every component of $f(\Sigma(f))$ bounds a region (actually, a $2$-disc) in $\mathbb{R}^{2}$. Given a component $\gamma$ of $f(\Sigma(f))$, if the number of points in the preimage of a point in $\mathbb{R}^{2}$ near $\gamma$ in the bounded region is greater than that of a nearby point in the unbounded region, then we say that $\gamma$ is positive; otherwise it is negative. Suppose $\Sigma(f)$ has $n$ negative components and $p$ positive components. We can perform surgeries of index $2$ along the $n$ negative components; in other words, for each such component, we remove an annular neighborhood of the corresponding singular point set component in $M$, and attach two $2$-discs on which we consider the natural embedding into $\mathbb{R}^{2}$. In this way, we can construct an image simple fold map of the resulting surface $M^{\prime}$ with $p$ positive components. Then $M^{\prime}$ is a disjoint union of $p$ copies of the $2$-sphere, and the Euler characteristic of $M^{\prime}$ is equal to $2p=\chi(M)+2n$. Thus, for the map $f$, the algebraic number of fold components is precisely equal to $\chi(M)/2$.
Let us now show that even if we allow $g$ not to be image simple, there exists no fold map $g\colon S^{2}\to\mathbb{R}^{2}$ with $|\Sigma(g)|$ even. Let $f\colon M\to\mathbb{R}^{2}$ be a fold map of a closed orientable surface $M$ to the plane. Then, $\Sigma(f)$ splits $M$ into compact orientable (possibly disconnected) surfaces $M_{+}$ and $M_{-}$ with boundary, where $f|_{M_{+}}$ (or $f|_{M_{-}}$) is orientation preserving (resp. reversing). By [6], their Euler characteristics coincide. Since $\chi(M_{+})+\chi(M_{-})=\chi(M)$ and $\chi(M)$ is even, we see that $\chi(M_{+})=\chi(M_{-})=\chi(M)/2$. Now, we have $\partial M_{+}=\partial M_{-}=\Sigma(f)$ and the number of boundary components of $M_{+}$ (or $M_{-}$) has the same parity as $\chi(M_{+})$ (resp. $\chi(M_{-})$), since they are orientable. Therefore, $|\Sigma(f)|$ has the same parity as $\chi(M)/2$, even if $f$ may not be image simple.
@@
We can construct a generic homotopy between the image simple fold maps $\tilde{w}$ and $\tilde{z}\colon S^{1}\tilde{\times}S^{2n-1}\to\mathcal{M}$ as follows. We first apply the birth move (see [20, Fig. 5]) to $\tilde{w}$ in such a way that we have a wrinkle consisting of two cusps together with fold arcs of indices n1n-1 and $n$ near the center circle of $\mathcal{M}$, i.e., inside the region $R$ bounded by $)\mathcal{M}\setminus\tilde{w}(\Sigma(\tilde{w}$)). Note that if we reverse the normal direction, the indices turn to $n$ and n1n-1. Then, we merge the two cusps using a path connecting them in the source manifold whose image is embedded in $R$. This is possible for $n>1$, since the fibers over the points in $R$ are connected. For $n=1$, the fibers are not connected: however, we can use a similar cusp elimination technique as in [15]. Then, we get an image simple fold map equivalent to $\tilde{z}$. The singular point set of the (generic) homotopy thus constructed is the disjoint union of a Möbius band and a cylinder, and it has no triple self-intersections.
@@ Acknowledgment
The authors would like to thank David Gay who kindly explained to the authors how to construct generalized Dehn twists. They would like to thank Sergey Melikhov for a discussion which led to Remark 6.3. They are also thankful to Liam Kahmeyer for discussions of the topic. They would also like to thank the anonymous referee for valuable comments which improved the presentation of the paper. The second author would also like to thank the Department of Mathematics, Kansas State University, for the hospitality during the preparation of the manuscript.